John has 2  marbles. Karen meets up with John on the way to Paul’s house. Karen has 5 marbles of her own. When they get to Paul’s house, he shows them his 9 marbles. What percentage of all of the marbles does John have? (Provide your answer to at least three digits.)


1
Expert's answer
2021-04-30T07:42:02-0400

Total number of marbles = 2+5+9= 16

John's percentage= 2/16 x 100= 12.5%